# Stong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

### v1

@@ Appendix A Additional proofs
$\leq c_{W_{q}}\delta^{q}\log^{2q}(\delta^{-1}){\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big{[}\mathds{1}_{\{f(z,\widetilde{z}(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{]}\Big{|}_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big{]}$
$=c_{W_{q}}\delta^{q}\log^{2q}(\delta^{-1}){\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big{]}$
@@
Hence,

### v2

@@ Appendix A Additional proofs
$\leq c_{W_{q}}\delta^{q}\log^{2q}(\delta^{-1}){\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big[\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big]\Big|_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big$]
$=c_{W_{q}}\delta^{q}\log^{2q}(\delta^{-1}){\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big]$
@@
Hence, for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,